what is logx - log (x+5) = -1? Thank you.
Answer by user_dude2008(1825)

all logs are base 10

log (x) - log (x+5) = -1


log[ (x)/(x+5) ] = -1


(x)/(x+5) = 10^(-1)

(x)/(x+5) = 1/10


10x=x+5

9x=5

x=5/9

Ans: x=5/9